OpenStudy (anonymous):

Solve using the quadratic formula. Round your answer to the nearest tenth: 3x2 = 5x + 7 A. {2.1, 0.4} B. {-2.1, 0.4} C. {-2.6, 0.9} D. {2.6, -0.9}

OpenStudy (anonymous):

First you need to know the quadratic formula, do you?

OpenStudy (anonymous):

no i dont .. i dont have my notes on me.

OpenStudy (anonymous):

well that is pretty important to know

OpenStudy (anonymous):

when you have: ax^2+bx+c=0 (where a, b and c are just numbers) Than the solution for this equation is: http://en.wikipedia.org/wiki/Quadratic_equation

OpenStudy (chaise):

The quadratic formula is: \[x = \frac{-b \pm \sqrt{b^{2}-4ac}}{2a}\] You will get two answers in your equation. a = the coefficient to the number you are squaring. b = the coefficient of the linear term c = the constant with no variable attached.

OpenStudy (chaise):

If all you're looking for is the answer - it is option D. If you need further help on Quadratics - I can help you with an in depth explanation and walk through of the question you are asking.

OpenStudy (anonymous):

okay...thank you!

OpenStudy (anonymous):

you are a hero chaise!

OpenStudy (chaise):

:) thanks sir. your accolades are appreciated.

OpenStudy (anonymous):

hahaha, anytime

OpenStudy (anonymous):

isn't the discriminant less than zero?